multiple choice Q

To what volume will a sample of gas expand if it is heated from 50.0 °C and 2.33 L to 500.0 °C?

17.9 L
9.74 L
4.38 L
10.3 L
5.58 L
Answer #1

fromcharles law

V1/T1 = V2/T2

V1 = 2.33 L

T1 = 50 C = 323 K

V2 = ?

T2 = 500+273.15 = 773.15 K

(2.33/323.15) = (V2/773.15)

V2 = 5.57 L

answer: 5.58 L